# Bine Trees: Enhancing Collective Operations by Optimizing Communication Locality

### v1

@@ 5.2.1. Comparison with Binomial Trees
The results show that Bine outperforms binomial trees in most scenarios. For half of the collectives, Bine is the best-performing algorithm in over 90% of the cases and matches the performance of binomial trees in the remaining ones.. On broadcast we see higher advantages than on LUMI, especially for vectors smaller than 128 KiB, where performance gain where of 4% on LUMI but of 42% on Leonardo. This can be attributed to MPICH relying on distance-halving broadcast (MPICH, 2025a), whereas Open MPI relies on distance-doubling broadcast (MPI, 2025a, b), which increases the number of bytes transmitted over global links, as we shown in Fig. 1.
@@ 5.2.2. Comparison with Other State-of-the-Art Algorithms
We summarize the results in Fig. 9(b). For allgather on $2048$ nodes, vectors smaller than 64 MiB benefit from the remapping approach described in Sec.4.4, achieving $2\times$ performance improvements compared to algorithms like Swing (De Sensi et al., 2024a) or sparbit (Loch and Koslovski, 2021; MPI, 2025d), which send non-contiguous data and introduce significant overhead. For reduce-scatter, Bine is the top-performing algorithm in 63% of the cases. For alltoall, results are similar to those on LUMI, with Bine being the best algorithm on small vectors. Finally, Bine achieves higher performance on broadcast compared to LUMI, consistently with our findings in Sec. 5.2.1.
@@
Figure 9. Comparison with state-of-the-art algorithms on Leonardo (N = binomial, D = default, R = ring).
[…]
Comparison with state-of-the-art algorithms on Leonardo (N = binomial, D = default, R = ring).
@@ 5.3. Evaluation on MareNostrum 5
MareNostrum 5 (Banchelli et al., 2025) uses a 2:1 oversubscribed fat-tree network topology based on InfiniBand NDR200. Each full-bandwidth sub-tree consists of 160 nodes. We utilized the ACC partition, where each node is equipped with $2\times$ Intel Sapphire Rapids 8460Y+ CPUs, 512 GB of DDR5 memory, and four 200 Gb/s NICs. Additionally, each node has 4x Nvidia Hopper GPUs with 64 GB of HBM2 memory.
@@ 5.3.2. Comparison with Other State-of-the-Art Algorithms
Fig. 10(a) summarizes the comparison between Bine and the best state-of-the-art algorithms. The results are consistent with our observations on the other platforms. For alltoall, gather, and scatter, Bine is the best-performing algorithm in fewer cases. This is primarily due to the smaller scale of the experiments (up to 64 nodes), compared to $1024$ or $256$ nodes on LUMI and Leonardo. At these scales, linear algorithms often outperform logarithmic ones, especially for large vector sizes. Conversely, for vectors smaller than 1 MiB, Bine outperformed the other algorithms by up to $2\times$ on 64 nodes for scatter.
@@
Figure 10. Comparison with state-of-the-art algorithms.
[…]
Comparison with state-of-the-art algorithms.
@@ 5.4. Evaluation on Fugaku
Fugaku (Sato et al., 2022) is based on the Tofu Interconnect D (Ajima et al., 2018) and uses a 6D toroidal network. In practice, each job can be allocated a 3D sub-torus. The system is composed of almost 159,000 nodes, each equipped with a Fujitsu A64FX CPU, 32 GiB of HBM memory, and six TNIs (i.e., NICs), each capable of injecting 54.4 Gb/s.
@@ 5.4.1. Bine Optimization for Torus Networks
So far, we have not made any assumptions about the network topology. However, we can extend Bine trees to better exploit communication locality on torus networks. Instead of considering a flat rank space, we treat ranks as coordinates in a multidimensional space, reducing distance between communicating ranks, as typical in algorithms optimized for toroidal networks (Jain and Sabharwal, 2010; De Sensi et al., 2024a). Since each node has six NICs, we divide the vector used in the collective into six parts. Then, we execute six collectives in parallel, each operating on a different part of the vector, using a separate NIC, and transmitting in a different direction of the 6D torus. To ensure that each collective uses a different NIC, we implemented the Bine algorithms directly on top of the low-level uTofu library (Fujitsu, 2025).
@@ 5.4.3. Comparison with Other State-of-the-Art Algorithms
As with other systems, we compare both with the default algorithm selected by Fujitsu MPI and with all other available algorithms by explicitly forcing their selection. We performed our analysis on 2x2x2, 4x4x4, 8x8x8, 64x64, and 32x256 nodes jobs. We report results in Fig. 10(b). For allreduce, Bine trees outperformed state-of-the-art algorithms in 62% of cases, especially for vectors larger than 2 KiB. The highest gains—up to $4.12\times$—were seen for 128 KiB–64 MiB vectors on $8192$ (32x256) nodes. While Fugaku uses the torus-optimized Trinaryx algorithm (Ajima et al., 2018; Hosono et al., 2021; Fujitsu, 2025), which builds three edge-disjoint spanning trees, its linear number of steps limits scalability compared to Bine’s logarithmic approach.
@@ 7. Related Work
Locality-Aware and Hierarchical Algorithm Locality-aware, and hierarchical or multi-leader algorithms (Bienz et al., 2022; Subramoni et al., 2011; Bayatpour et al., 2017; Hidayetoglu et al., 2024) have been proposed for various collectives. These algorithms, however, either require knowledge of the underlying topology (Subramoni et al., 2011) or make assumptions about the partitioning of ranks across groups (Bienz et al., 2022). Unlike hierarchical methods, Bine trees are as generic as traditional collective algorithms and, as shown in Sec. 5.3.2, can also be effectively integrated into hierarchical algorithms.
Automatic Algorithm Synthesis Some methods aim to automatically determine the optimal collective algorithm based on network topology and collective specifications (Shah et al., 2022; Cai et al., 2021; Won et al., 2024). However, their complexity increases exponentially as the system scales. For example, computing a solution for 128 nodes can take up to 11 hours (Shah et al., 2022), and a new solution may need to be recalculated for each allocation, which makes these approaches not practical on shared supercomputers, where users do not have control over allocations.
Topology-Optimized Algorithms Researchers developed collective algorithms tailored to specific network topologies (Almási et al., 2005; Ma et al., 2021; Feng et al., 2022). In contrast, Bine trees are versatile and topology-agnostic and consistently enhance performance across a variety of topologies.
@@ Acknowledgements.
We thank Daniele Di Bari (CINECA) for supporting us in executing large-scale tests on Leonardo, Kim McMahon (HPE Cray) for the support with Cray MPICH collective algorithms selection, Kurt Lust for the information on mapping of nodes to groups on LUMI, and Faveo Hörold (ETH Zürich and RIKEN) for helping with preliminary tests on Fugaku. This work is supported by the European Union Commission’s Horizon Europe, under grant agreements 101175702 (NET4EXA), by the Sapienza University Grants ADAGIO and D2QNeT (Bando per la ricerca di Ateneo 2023 and 2024), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement PSAP, No. 101002047). Daniele De Sensi is a member of Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). The authors used ChatGPT to assist with editing. All ideas, content, and conclusions are the authors’ own.

### v2

@@ 5.2.1. Comparison with Binomial Trees
The results show that Bine outperforms binomial trees in most scenarios. For half of the collectives, Bine is the best-performing algorithm in over 90% of the cases and matches the performance of binomial trees in the remaining ones.. On broadcast we see higher advantages than on LUMI, especially for vectors smaller than 128 KiB, where performance gain where of 4% on LUMI but of 42% on Leonardo. This can be attributed to MPICH relying on distance-halving broadcast (mpich-bcast-binomial), whereas Open MPI relies on distance-doubling broadcast (ompi-bcast-binomial; ompi-binomial), which increases the number of bytes transmitted over global links, as we shown in Fig. 1.
@@ 5.2.2. Comparison with Other State-of-the-Art Algorithms
We summarize the results in Fig. 10(b). For allgather on $2048$ nodes, vectors smaller than 64 MiB benefit from the remapping approach described in Sec.4.4, achieving $2\times$ performance improvements compared to algorithms like Swing (swing) or sparbit (sparbit; ompi-allgather-sparbit), which send non-contiguous data and introduce significant overhead. For reduce-scatter, Bine is the top-performing algorithm in 63% of the cases. For alltoall, results are similar to those on LUMI, with Bine being the best algorithm on small vectors. Finally, Bine achieves higher performance on broadcast compared to LUMI, consistently with our findings in Sec. 5.2.1.
@@
Figure 10. Comparison with state-of-the-art algorithms on Leonardo (N = binomial, D = default, R = ring). :
@@ 5.3. Evaluation on MareNostrum 5
MareNostrum 5 (marenostrum) uses a 2:1 oversubscribed fat-tree network topology based on InfiniBand NDR200. Each full-bandwidth sub-tree consists of 160 nodes. We utilized the ACC partition, where each node is equipped with $2\times$ Intel Sapphire Rapids 8460Y+ CPUs, 512 GB of DDR5 memory, and four 200 Gb/s NICs. Additionally, each node has 4x Nvidia Hopper GPUs with 64 GB of HBM2 memory.
@@ 5.3.2. Comparison with Other State-of-the-Art Algorithms
Fig. 11(a) summarizes the comparison between Bine and the best state-of-the-art algorithms. The results are consistent with our observations on the other platforms. For alltoall, gather, and scatter, Bine is the best-performing algorithm in fewer cases. This is primarily due to the smaller scale of the experiments (up to 64 nodes), compared to $1024$ or $256$ nodes on LUMI and Leonardo. At these scales, linear algorithms often outperform logarithmic ones, especially for large vectors. Conversely, for vectors smaller than 1 MiB, Bine outperformed the other algorithms by up to $2\times$ on 64 nodes.
@@
Figure 11. Comparison with state-of-the-art algorithms. :
@@ 5.4. Evaluation on Fugaku
Fugaku (fugaku) is based on the Tofu Interconnect D (tofu) and uses a 6D toroidal network. In practice, each job can be allocated a 3D sub-torus. The system is composed of almost 159,000 nodes, each equipped with a Fujitsu A64FX CPU, 32 GiB of HBM memory, and six TNIs (i.e., NICs), each capable of injecting 54.4 Gb/s.
@@ 5.4.1. Bine Optimization for Torus Networks
So far, we have not made any assumptions about the network topology. However, we can extend Bine trees to better exploit communication locality on torus networks. Instead of considering a flat rank space, we treat ranks as coordinates in a multidimensional space, reducing distance between communicating ranks, as typical in algorithms optimized for toroidal networks (bucket1; swing). Since each node has six NICs, we divide the vector used in the collective into six parts. Then, we execute six collectives in parallel, each operating on a different part of the vector, using a separate NIC, and transmitting in a different direction of the 6D torus. To ensure that each collective uses a different NIC, we implemented the Bine algorithms directly on top of the low-level uTofu library (utofu). We provide a more detailed description of the design and the implementation of Bine trees for toroidal networks in Appendix D.
@@ 5.4.3. Comparison with Other State-of-the-Art Algorithms
As with other systems, we compare both with the default algorithm selected by Fujitsu MPI and with all other available algorithms by explicitly forcing their selection. We performed our analysis on 2x2x2, 4x4x4, 8x8x8, 64x64, and 32x256 nodes jobs. We report results in Fig. 11(b). For allreduce, Bine trees outperformed state-of-the-art algorithms in 62% of cases, especially for vectors larger than 2 KiB. The highest gains—up to $4.12\times$—were seen for 128 KiB–64 MiB vectors on $8192$ (32x256) nodes. While Fugaku uses the torus-optimized Trinaryx algorithm (tofu; oai:ipsj.ixsq.nii.ac.jp:00214220; topotofu), which builds three edge-disjoint spanning trees, its linear number of steps limits scalability compared to Bine’s logarithmic approach.
@@ 7. Related Work
Locality-Aware and Hierarchical Algorithm Sevaral locality-aware, and hierarchical or multi-leader algorithms (10.1145/3555819.3555825; 6061150; mleader; hidayetoglu2024hicclhierarchicalcollectivecommunication) have been proposed. These algorithms, however, either require knowledge of the underlying topology (6061150) or make assumptions about the partitioning of ranks across groups (10.1145/3555819.3555825). On the other hand, Bine trees are as generic as binomial trees and, as shown in Sec. 6, can be effectively integrated into hierarchical algorithms.
Automatic Algorithm Synthesis Some methods aim to automatically determine the optimal collective algorithm based on network topology and collective specifications (shah2022taccl; 10.1145/3437801.3441620; won2024tacostopologyawarecollectivealgorithm). However, their complexity increases exponentially as the system scales. For example, computing a solution for 128 nodes can take up to 11 hours (shah2022taccl), and a new solution may need to be recalculated for each allocation, which makes these approaches not practical on shared supercomputers, where users do not have control over allocations.
Topology-Optimized Algorithms Researchers developed collective algorithms tailored to specific network topologies (10.1145/1088149.1088183; 9644896; 10.1145/3524059.3532380). In contrast, Bine trees are versatile and topology-agnostic and consistently enhance performance across a variety of topologies.
@@ Acknowledgements.
We thank Daniele Di Bari (CINECA) for support with large-scale tests on Leonardo, Kim McMahon (HPE Cray) for assistance with Cray MPICH collective algorithm selection, Kurt Lust for information on node-to-group mapping on LUMI, and Faveo Hörold (ETH Zürich and RIKEN) for help with preliminary tests on Fugaku. This work is supported by the European Union’s Horizon Europe under grant 101175702 (NET4EXA), by Sapienza University Grants ADAGIO and D2QNeT (Bando per la ricerca di Ateneo 2023 and 2024), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant PSAP, No. 101002047). Daniele De Sensi is a member of GNCS-INdAM. The authors used ChatGPT for editing; all ideas, content, and conclusions are their own.
Appendix A Detailed Description of Distance-Doubling Bine Trees
Since we want to preserve the $33\%$ modular distance reduction compared to distance-doubling binomial trees, in distance-doubling Bine trees, ranks communicate in the opposite order compared to distance-halving butterflies. Specifically, after receiving data at step $i$ (as described in Sec. 3.2), at each subsequent step $j>i$ a rank $r$ sends data to a rank $q$ determined as:
(5) $q=\begin{cases}\Big(r+\sum_{k=0}^{j}(-2)^{k}\Big)\bmod p,&\text{if }r\text{ is even},\\ \Big(r-\sum_{k=0}^{j}(-2)^{k}\Big)\bmod p,&\text{if }r\text{ is odd}.\end{cases}$
The modular distance between two communicating ranks is thus $\sum_{k=0}^{j}(-2)^{k}$, implying a $33\%$ reduction in distance compared to standard distance-doubling trees (for the reasons discussed in Sec. 2.4.1). Note that $\sum_{k=0}^{j}(-2)^{k}=\tfrac{1}{3}-\tfrac{-2^{j+1}}{3}$ is always odd, meaning that even ranks always communicate directly with odd ranks, and vice versa.
We now analyze when the data sent by the root rank $r$ reaches a given rank $q$. For simplicity, we consider a broadcast with root $r=0$, but the same reasoning applies to other collectives and to arbitrary roots $r\neq 0$ (in that case, the tree is simply rotated by $r$ positions). At step $j$, the root $r=0$ sends data to rank
$q=\sum_{k=0}^{j}(-2)^{k}\bmod p.$
$q$ is odd, and when $q$ later forwards the data at some step $c>j$, it sends it to:
$t=q-\sum_{k=0}^{c}(-2)^{k}\bmod p=\sum_{k=0}^{j}(-2)^{k}-\sum_{k=0}^{c}(-2)^{k}\bmod p.$
i.e., the sign between each summation alternates from positive to negative and vice-versa.
To generalize, consider the path through which data from root 0 reaches a rank $q$ after $h$ steps $s_{0}<s_{1}<\dots<s_{h-1}$. If $q$ is odd, since communication alternates between even and odd ranks, $h$ must be odd, and:
$q=\sum_{i=0}^{s_{0}}(-2)^{i}-\sum_{i=0}^{s_{1}}(-2)^{i}+\sum_{i=0}^{s_{2}}(-2)^{i}-\dots+\sum_{i=0}^{s_{h-1}}(-2)^{i}\mod p.$
Simplifying the summations gives:
(6) $q=\sum_{i=0}^{s_{0}}(-2)^{i}+\sum_{i=s_{1}+1}^{s_{2}}(-2)^{i}+\dots+\sum_{i=s_{h-2}+1}^{s_{h-1}}(-2)^{i}\mod p.$
The right-hand side corresponds to a negabinary representation, whose least significant bit is $1$, as illustrated in Fig. 12 for a 12-bit string with $s_{0}=1$, $s_{1}=5$, and $s_{2}=8$.
$s_{2}$ $s_{1}$ $s_{0}$
[…]
0 0 0 1 1 1 0 0 0 0 1 1
Figure 12. Negabinary representation corresponding to Eq. 6 with $s_{0}=1$, $s_{1}=5$, and $s_{2}=8$.
In this representation, the leftmost and rightmost positions of each contiguous sequence of $1$s correspond to the values of the $s_{i}$. Thus, if $q$ is odd and its negabinary encoding matches Fig. 12, the data from the root reached $q$ through step $s_{0}=1$, then $s_{1}=5$, and finally $s_{2}=8$.
To isolate the positions of the $s_{i}$ (i.e., to create a binary string with 1s only at those positions, e.g., $000100100010$ for the example just described), we define $h(q,p)$ as the negabinary string in Fig. 12, and compute:
$\nu(q,p)=h(q,p)\oplus(h(q,p)\gg 1).$
In essence, by inspecting the bits of $\nu(q,p)$, we can reconstruct exactly how the data propagated from the root to rank $q$.
Rank $q$ receives the data at step $s_{h-1}$, corresponding to the most significant bit set to 11 in $\nu(q,p)$. At each subsequent step $j$, it forwards the data to a rank whose $\nu$ differs only in bit jj—that is, the destination rank has a 11 in position $j$ and 0s in all more significant positions, consistent with the algorithm’s design.
The same reasoning extends to even ranks. In that case, $h$ is even, and:
$q=\sum_{i=0}^{s_{0}}(-2)^{i}-\sum_{i=0}^{s_{1}}(-2)^{i}+\sum_{i=0}^{s_{2}}(-2)^{i}-\dots-\sum_{i=0}^{s_{h-1}}(-2)^{i}\mod p,$
which simplifies to:
$q=-\sum_{i=s_{0}+1}^{s_{1}}(-2)^{i}-\sum_{i=s_{2}+1}^{s_{3}}(-2)^{i}-\dots-\sum_{i=s_{h-2}+1}^{s_{h-1}}(-2)^{i}\mod p.$
Multiplying both sides by 1-1 yields:
$q=\sum_{i=s_{0}+1}^{s_{1}}(-2)^{i}+\sum_{i=s_{2}+1}^{s_{3}}(-2)^{i}+\dots+\sum_{i=s_{h-2}+1}^{s_{h-1}}(-2)^{i}\mod p.$
Adding $p$ to both sides, we obtain:
(7) $p-q=\sum_{i=s_{0}+1}^{s_{1}}(-2)^{i}+\sum_{i=s_{2}+1}^{s_{3}}(-2)^{i}+\dots+\sum_{i=s_{h-2}+1}^{s_{h-1}}(-2)^{i}\mod p.$
The right-hand side again represents a negabinary number, but with its least significant bit equal to 0, as shown in Fig. 13 for $s_{0}=1$, $s_{1}=5$, $s_{2}=8$, and $s_{3}=10$. For this reason, for even $q$ we define:
$\nu(q,p)=h(p-q,p)\oplus(h(p-q,p)\gg 1).$
$s_{3}$ $s_{2}$ $s_{1}$ $s_{0}$
[…]
0 1 1 0 0 0 1 1 1 1 0 0
Figure 13. Negabinary representation corresponding to Eq. 7, with $s_{0}=1$, $s_{1}=5$, $s_{2}=8$, and $s_{3}=10$.
Appendix B Comparison of Non-Contiguous Data Transfers Techniques
Figure 14. Comparison of techniques for handling the transmission of non-contiguous data (see Sec. 4.3.1). Results are obtained on LUMI for an allgather collective operation. Each cell shows a letter corresponding to the best technique (B=Block-by-block, P=Permute, S=Send, T=Two Transmissions), and the performance gain compared to standard binomial butterflies.:
[…]
For larger vectors, the cost of permutation further increases because more data must be rearranged, making the block-by-block technique more efficient. However, as the number of nodes grows, the number of transmissions also rises, and the best-performing approach becomes the two transmissions technique. Finally, we note that the current implementation still offers opportunities for optimization. For instance, in the permute variant, data permutation could be overlapped with communication by employing an additional buffer, in which data received during the previous step is permuted while data for the current step is transmitted.
Appendix C Managing Non-Power of Two Ranks
If the number of ranks is not a power of $2$, different strategies can be adopted. The most straightforward approach follows techniques commonly used in binomial trees (mpich-bcast-binomial). For clarity, consider a one-dimensional torus of size $p$ not equal to a power of two. Let $p^{\prime}=2^{\lfloor\log_{2}p\rfloor}$. For tree-based collectives (e.g., a broadcast), one option is to first execute the collective among $p^{\prime}$ ranks, and then have the remaining $p-p^{\prime}$ ranks receive the data from $p-p^{\prime}$ of the already-participating ranks.
For butterfly-based collectives (e.g., reduce-scatter), a straightforward solution is to have the last $p-p^{\prime}$ ranks forward their data to the first $p-p^{\prime}$ ranks, then execute the reduce-scatter on these $p^{\prime}$ ranks, and finally redistribute part of the resulting data back to the last $p-p^{\prime}$ ranks. However, this doubles the total communication volume, since the first $p-p^{\prime}$ ranks must first receive an entire vector from one of the last ranks before participating in the reduce-scatter.
Figure 15. Bine trees constructed for a number of ranks that is not a power of $2$.
If $p$ is even, we can adopt a simpler approach that does not increase the total communication volume. In this case, the collective is executed as usual, but some ranks may be reached twice. This situation is illustrated in Fig. 15 for collectives with $6$ and $10$ nodes. For simplicity, a tree is shown, but the same applies to a butterfly, which, as discussed in Sec. 3.1, can be seen as the superposition of multiple trees, each rooted at a different rank.
For example, in the $6$-node case, both rank $4$ and rank $5$ are reached twice. This would compromise correctness. For instance, in a reduce operation, the same contribution might be applied twice. To avoid this issue, it suffices to remove one of the two subtrees rooted at the rank that is reached twice. For example, for rank $4$ we can either remove it as a leaf (and the children of rank $5$), or remove the subtree rooted at rank $4$ (which has rank $5$ as a child).
To determine which of the two subtrees to remove, let us consider the $10$-node tree shown in Fig. 15. Each node is annotated with the step at which the data from the root reaches it (e.g., node $8$ is reached at step $3$, since node 0 sends the data to node $9$ at step $2$, and then node $9$ forwards it to node $8$ at step $3$). Since the subtree reached later always contains fewer nodes (which are also contained in the other subtree), we discard that one, as removing the larger subtree could compromise correctness. This is evident in the case of node $5$. We remove the subtree reached at step $3$ (Fig. 15 [Uncaptioned image]), since removing the other one (Fig. 15 [Uncaptioned image]) would disconnect a larger portion of the tree, leaving some nodes unreachable (e.g., nodes $6$ and $7$).
It is worth noting that this approach cannot be directly applied if $p$ is odd. Since communication always occurs in pairs, some ranks would inevitably attempt to communicate with a rank that is already engaged. In such cases, one must revert to the standard techniques discussed above.
Figure 16. Optimization of a distance-halving Bine tree for a 2D torus. For both Bine tree and torus-optimized Bine tree we show both the 2D torus on which the tree is built, and the corresponding tree.
Appendix D Optimization for Torus Networks
We now describe how to extend Bine trees to better exploit communication locality on torus networks. For simplicity, we focus on trees; however, as discussed in Sec. 3.1, a butterfly can be viewed as a superposition of multiple trees, each rooted at a different rank. Therefore, the considerations in this section apply to both binomial trees and butterflies. We first assume that each torus dimension is a power of $2$, and then generalize to arbitrary dimension sizes.
D.1. Selection of Communicating Partner
We illustrate this in Fig. 16, comparing a standard Bine tree with its torus-optimized counterpart on a $4\times 4$ torus. Standard Bine trees optimize for modular distance assuming a one-dimensional torus, which does not necessarily yield efficient communication paths in higher-dimensional topologies. For example, in the left part of Fig. 16, edge labels show the modular distance between communicating ranks as if they were arranged on a 1D torus. A sign indicates communication direction: a minus sign denotes communication to the left of the source rank. For instance, the modular distance between ranks 0 and $15$ is computed as:
$d(0,15)=\min((0-15)\bmod 16,\ (15-0)\bmod 16)=\min(1,15)=1,$
which we label as 1-1 since the shorter path wraps leftward. However, on a 2D torus, the actual distance between ranks 0 and $15$ is 2 hops (see Fig. 16 [Uncaptioned image]), traversing multiple links and potentially introducing additional congestion, as discussed in Sec. 1.
To address this, we treat ranks as coordinates in a multidimensional space and apply the Bine construction dimension by dimension, similar to other torus-optimized algorithms (bucket1; swing). Communication partners are then chosen along single dimensions: instead of distances 5,3,1,1-5,3,-$1,$ rank 0 exchanges with ranks at (0,1),(1,0),(0,1),(1,0)(0,-1),(-1,0),(0,1),(1,0). Concretely, rank 0 at coordinates $(0,0)$ first communicates with $(0,3)$ (rank 3), then with $(3,0)$ (rank 12), then with $(0,1)$ (rank 1), and finally with $(1,0)$ (rank 4). As shown in Fig. 16 [Uncaptioned image], this approach reduces the number of crossed links, alleviating congestion and improving performance.
D.2. Data Handling
However, while this approach reduces communication distance, it complicates data management. To illustrate this issue and possible solutions, we consider a scatter operation implemented with a Bine tree. Other collectives are handled similarly; for example, in a gather, data flows from the leaves to the root, and in an allgather, each rank uses its own tree. The root splits the data into “blocks”, one per rank. Each node must forward to its children all the blocks belonging to the subtree rooted at that child. In standard distance-halving Bine trees, the descendants of any node are contiguous. For example, in Fig. 16, the subtree rooted at rank $11$ (highlighted by the dashed box) contains only consecutive ranks. As a result, when rank 0 sends data to rank $11$, it transfers a set of contiguous memory blocks (i.e., the blocks in the range $[6,13]$ Fig. 16 [Uncaptioned image]).
In contrast, in torus-optimized Bine trees the descendants of a node are no longer contiguous, since ranks are arranged according to multidimensional coordinates. Returning to the scatter example, in the first step rank 0 (with coordinates $(0,0)$) sends data to rank $3$ (with coordinates $(0,3)$). In this case, although those ranks are contiguous in the 2D torus (Fig. 16 [Uncaptioned image]), the blocks in the one-dimensional input vector are not (Fig. 16 [Uncaptioned image]). This increases data-handling complexity compared to the standard case, and can significantly hinder performance, especially for small- to medium-sized vectors (sparbit), as we show in Sec. 5.
An alternative is to permute the vector before transmission, ensuring that all subsequent sends by any rank involve contiguous data. This can be achieved by renumbering the nodes according to a DFS postorder traversal (i.e., the order obtained by a depth-first search where a node is visited after all its children). For instance, in Fig. 16, the block originally at position $10$ (coordinates $(2,2)$) is moved to position 0, the block at position $14$ (coordinates $(3,2)$) to position $1$, and so forth.
To avoid the initial permutation, ranks could send contiguous data as if the permutation had occurred. At the end of the collective, however, each rank would hold the wrong block; for instance, rank $10$ (coordinates $(2,2)$) would receive the data intended for rank 0 (coordinates $(0,0)$). This can be corrected by adding an additional communication step at the end, where rank $10$ sends its data to the correct rank (i.e., rank 0), and other ranks perform analogous exchanges.
Finally, some collectives are composed of multiple operations, such as an allreduce implemented as a reduce-scatter followed by an allgather. In these cases, even if the reduce-scatter places blocks in the wrong order, the subsequent allgather, which effectively inverts the reduce-scatter schedule, restores the blocks to the correct order.
D.3. Non-Power of Two
If one (or more) torus dimensions is not a power of $2$, we can rely on techniques commonly used in standard binomial trees and butterflies, as discussed in Appendix C. On a multidimensional torus, however, beyond the challenges already mentioned, such techniques complicate the selection of the $p^{\prime}$ ranks, particularly when trying to preserve short communication distances among participating ranks. As a consequence, several communication steps may require traversing long paths across multiple links.
Moreover, running the collective on a subtorus with $p^{\prime}$ ranks (whose dimensions are all powers of two) would disrupt the direct correspondence between torus coordinates and linear rank numbering, thereby complicating both process mapping and data management. When all torus dimensions are even, we instead adopt the approach described in Appendix C, which eliminates duplicated subtrees. This strategy, illustrated in Fig. 17 for a $2\times 6$ torus, does not require identifying a subtorus and does not add extra complexity compared to the power-of-two case.
Figure 17. Torus-optimized Bine trees for toroidal networks with dimensions that are not powers of $2$.
D.4. Multi-ported Torus
On many torus networks, nodes can transmit simultaneously in all directions. For example, on Fugaku each node is equipped with six TNIs (i.e., NICs), each capable of injecting $54.4$ Gb/s. To fully exploit this capability, on a $D$-dimensional torus we divide the vector used in the collective into $2D$ parts. We then execute $2D$ collectives in parallel, each operating on $n/2D$ bytes, using a separate NIC and transmitting in a different direction.
For instance, as shown in Fig. 18 for a $2D$ torus, rank 0 would normally communicate sequentially in the order east (E) $\rightarrow$ north (N) $\rightarrow$ west (W) $\rightarrow$ south (S). With port parallelization, this order is applied only to $1/4$ of the data. In parallel, another $1/4$ is transmitted in the order N $\rightarrow$ W $\rightarrow$ S $\rightarrow$ E, another $1/4$ with order W $\rightarrow$ S $\rightarrow$ E $\rightarrow$ N, and the last $1/4$ with order S $\rightarrow$ E $\rightarrow$ N $\rightarrow$ W. In this way, all four NICs are active in parallel, thus saturating the available bandwidth. It is worth remarking that on rectangular torus, after some step some of the ports will not be used anymore, since no more communications will happen along those shorter dimensions.
Figure 18. Communications happening on a 2D torus when multiported Bine tree algorithms are used.
To ensure that each collective uses a different NIC, we implemented the Bine algorithms directly on top of the low-level uTofu library, caching registered remote addresses as we suppose it would likely happen in Fujitsu MPI as well. To simplify the management of several interfaces, each one is managed by a separate thread.